OpenStudy (anonymous):

Choose the replacement for x and y that makes this equation a true statement: 2x + 55y = −214

OpenStudy (anonymous):

2x + 55y = -214 2(-24.5) + 55(-3) = -214 -49 - 165 = -214 x = -24.5 y = -3 (I think so anyways.)

OpenStudy (anonymous):

there is an infinite number of solutions to this equation. Choose any number for x (which we'll call t) and rearrange the equation to isolate y (y=(-214-2x)/55) then y will be given from inserting t into the equation.

OpenStudy (anonymous):

plotting the graph will give you all the possible combinations of x and y.